### v2

# Mutual Information Optimization for SIM-Based Holographic MIMO Systems

Abstract
In the context of emerging stacked intelligent metasurface (SIM)-based holographic MIMO (HMIMO) systems, a fundamental problem is to study the mutual information (MI) between transmitted and received signals to establish their capacity. However, direct optimization or analytical evaluation of the MI, particularly for discrete signaling, is often intractable. To address this challenge, we adopt the channel cutoff rate (CR) as an alternative optimization metric for the MI maximization. In this regard, we propose an alternating projected gradient method (APGM), which optimizes the CR of a SIM-based HMIMO system by adjusting signal precoding and the phase shifts across the transmit and receive SIMs on a layer-by-layer basis. Simulation results indicate that the proposed algorithm significantly enhances the CR, achieving substantial gains, compared to the case with random SIM phase shifts, that are proportional to those observed for the corresponding MI. This justifies the effectiveness of using the channel CR for the MI optimization. Moreover, we demonstrate that the integration of digital precoding, even on a modest scale, has a significant impact on the ultimate performance of SIM-aided systems.
Index Terms:
Channel CR, MI, SIM, HMIMO, optimization.
I Introduction
Intelligent metasurfaces are expected to significantly impact the evolution of future wireless communications. These surfaces, composed of a large number of controllable metamaterial elements, can dynamically modify the electromagnetic (EM) wave propagation [1]. This ability enables them to improve energy efficiency [2], network capacity, reliability [3] and connectivity [4], while also supporting other functionalities, such as localization and sensing. However, their advantages hinge on acquiring accurate channel state information (CSI), a challenging task for conventional structures, such as reconfigurable intelligent surfaces. Furthermore, the path loss in the RIS-assisted links can substantially diminish these benefits, indicating that placing intelligent metasurfaces close to transmitters and receivers maximizes their potential.
The above discussions naturally motivate the integration of intelligent metasurface structures with wireless communication transceivers in massive MIMO (mMIMO) and HMIMO systems [5]. The use of intelligent metasurfaces in these systems offers a tradeoff between minimizing the number of RF chains and increasing control over EM fields [6]. Moreover, they can provide beamforming capabilities comparable to those of conventional phased array antennas, but at lower cost and power consumption. However, limitations due to their typically single-layer structure and the finite number of tunable states of their elements in practice can result in beam misalignment, potentially compromising the expected performance.
To address the above limitations, multi-layer metasurface structures have emerged as a promising solution for flexibly forming different radiation patterns compared to their single-layer counterparts [7]. Such structures, called SIMs, were recently introduced [8], and their origin was inspired from the architecture of a deep neural network [9]. Indeed, SIMs mirror the multi-layer structure of neural networks that are capable of modeling various functions with improved signal processing capabilities. Similarly, when metasurface layers are properly placed, SIM can implement signal processing directly in the EM wave domain. This approach can potentially reduce the reliance on digital beamforming and high-precision digital-to-analog converters.
In [10], the signal processing capabilities of SIMs were exploited for the implementation of 2D discrete Fourier transform (DFT) for direction of arrival (DOA) estimation. Moreover, a hybrid channel estimator, in which the received training symbols were processed first in the wave domain and later in the digital domain, was proposed in [11]. In [12], the authors studied the achievable sum-rate maximization for a downlink channel between a SIM-assisted base station and multiple single-antenna users. Integrating SIMs with transmitters and receivers into the so-called SIM-based HMIMO system, which performs signal precoding and combining in the wave domain, was proposed in [13]. Moreover, the achievable rate optimization for the SIM-based HMIMO system was studied in [14].
Despite quite a few papers about SIM-based systems, none of them has considered the achievable rate of such systems using discrete signaling, i.e., the MI. This gap has motivated us to investigate the optimization of the MI in the aforementioned systems. As shall be seen shortly, the direct optimization of the MI is intractable, and thus we instead propose the use of the channel CR as an alternative metric to facilitate the MI optimization [15, 16], which is justified as follows. First, the MI is related to the CR, $R_{0}$, by $\text{MI}\geq N_{r}(1-\log_{2}e)+R_{0}$ [15, Eqn. (36)], meaning that maximizing the CR accordingly translates into maximizing the MI. Second, $R_{0}$ and the codeword error probability, $P_{e}$, are related by $P_{e}\leq 2^{-n(R_{0}-R)}$ [17, Eq. (6.8–19)], where $n$ is the codeword length and $R$ is the information rate. Essentially, letting $n\rightarrow\infty$, $P_{e}$ can be made arbitrarily small provided that $R<R_{0}$. Since the channel capacity represents the theoretical maximum information rate for reliable communications, the CR is effectively employed as a practical lower bound on the channel capacity. The inter-connection among the CR, the MI, and the channel capacity indeed suggests that the CR can be used to optimize the MI in an indirect, yet efficient, way.
The main contributions of this paper are listed as follows:
  • We show that using the channel CR as an alternative optimization metric enables efficient maximization of the MI in SIM-based HMIMO systems. Specifically, the CR is expressed as a closed-form expression, which facilitates the derivation of an efficient optimization algorithm.
  • To maximize the CR, we formulate a joint optimization problem of transmit precoding and the phase shifts for both the transmit and the receive SIMs. An APGM is proposed to solve this problem, which particularly optimizes the phase shifts for the transmit and the receive SIMs on a layer-by-layer basis, using an independent step size for each layer, which is different from [13]. Also, we provide the computational complexity of the proposed algorithm.
  • We demonstrate through simulation results that the proposed algorithm can significantly increase the CR and the MI in a proportional manner. Also, the increase of the modulation order without changing the transmit signal power has a negligible influence on the MI. Finally, incorporating even a small scale digital precoding into a SIM-based HMIMO system leads to substantial improvements on the achievable MI.
Notation: Bold lower and upper case letters represent vectors and matrices, respectively. $\operatorname{Tr}(\cdot)$ denotes the trace of a matrix and $\text{diag}(\cdot)$ transforms a vector into a diagonal matrix. $\log_{2}(\cdot)$ is the binary logarithm, $\exp(\cdot)$ denotes the exponent, $\mathbb{E}\{\cdot\}$ denotes the expectation operation and $(\cdot)^{*}$ is the complex conjugate. $(\cdot)^{T}$ and $(\cdot)^{H}$ represent transpose and Hermitian transpose, receptively. $\nabla_{x}f(\cdot)$ is the complex gradient of $f(\cdot)$ with respect to (w.r.t.) $x^{*}$.
II System Model
We consider a SIM-based HMIMO communication system with $N_{t}$ transmit antennas and $N_{r}$ receive antennas, where both the transmitter and the receiver are equipped with SIMs. The transmit SIM consists of $L$ metasurface layers with $N$ meta-atoms in each layer, while the receive SIM consists of $K$ metasurface layers with $E$ meta-atoms in each layer. We assume that the signal propagation through every SIM layer is ideal, i.e., without any power loss. The SIM meta-atoms are connected to smart controllers, which can independently adjust the phase shift of each meta-atom. In this way, MIMO transceivers with closely spaced SIMs can implement signal beamforming directly in the EM wave domain.
For the $l$-th layer of the transmit SIM, the phase shift matrix is denoted as $\boldsymbol{\mathbf{\Phi}}^{l}=\text{diag}(\boldsymbol{\phi}^{l})=\text{diag}([\phi_{1}^{l}\>\phi_{2}^{l}\>\cdots\>\phi_{N}^{l}]^{T})$. Also, $\phi_{n}^{l}=\exp(j\theta_{n}^{l})$, where $\theta_{n}^{l}$ is the phase shift of the $n$-th meta-atom in the $l$-th layer. Signal propagation between the $(l-1)$-th and $l$-th layer of the transmit SIM is modeled by the matrix $\mathbf{W}^{l}\in\mathbb{C}^{N\times N}$, where $[\mathbf{W}^{l}]_{m,n}$ denotes the signal propagation between the $n$-th meta-atom of the $(l-1)$-th layer and the $m$-th meta-atom of the $l$-th layer. According to Rayleigh-Sommerfeld diffraction theory, $[\mathbf{W}^{l}]_{m,n}$ is equal [18, Eq. (1)]
$[\mathbf{W}^{l}]_{m,n}=\frac{A\cos\chi_{m,n}}{d_{m,n}}\Bigl{(}\frac{1}{2\pi d_{m,n}}-\frac{j}{\lambda}\Bigr{)}e^{j\frac{2\pi d_{m,n}}{\lambda}}$ (1)
for $l=2,3,\dots,L$, where $A$ is the area of each meta-atom, $\lambda$ is the wavelength, $d_{m,n}$ is the propagation distance between the meta-atoms in the $(l-1)$-th and $l$-th layer of the transmit SIM, $\chi_{m,n}$ is the angle between the propagation direction and normal direction of the $(l-1)$-th layer [^1] In this paper, it is assumed that all layers of the transmit SIM have the same arrangement and thus the right-hand side (RHS) of (1) does not depend on $l$. The same applies to the receive SIM.. Also, signal propagation between the transmit antenna array and the first layer of the transmit SIM is modeled by $\mathbf{W}^{1}\in\mathbb{C}^{N\times N_{t}}$, which is defined as (1). The wave domain beamforming matrix of the transmit SIM is given as $\mathbf{B}=\boldsymbol{\mathbf{\Phi}}^{L}\mathbf{W}^{L}\cdots\boldsymbol{\mathbf{\Phi}}^{2}\mathbf{W}^{2}\boldsymbol{\mathbf{\Phi}}^{1}\mathbf{W}^{1}\in\mathbb{C}^{N\times N_{t}}.$
[…]
Let $\mathbf{G}\in\mathbb{C}^{N\times E}$ be the channel between the transmit and the receive SIM. Then the end-to-end channel matrix for the considered communication system is $\mathbf{H}=\mathbf{Z}\mathbf{G}\mathbf{B}\in\mathbb{C}^{N_{r}\times N_{t}}$ and the signal vector at the receive antennas is given by
$\mathbf{y}=\mathbf{H}\mathbf{P}\mathbf{x}_{i}+\mathbf{n}$ (2)
where $\mathbf{P}\in\mathbb{C}^{N_{t}\times N_{s}}$ is the precoding matrix that satisfies the following the average transmit power constraint:
$\operatorname{Tr}(\mathbf{P}\mathbf{P}^{H})=N_{s}$ (3)
where $N_{s}\leq\min(N_{t},N_{r})$ is the number of the transmitted modulation symbols. The transmit vector $\mathbf{x}_{i}\in\mathbb{C}^{N_{s}\times 1}$ consists of elements chosen from a discrete symbol alphabet of size $M$, and thus, the number of different transmit vectors is $N_{vec}=M^{N_{s}}$. In addition, it is assumed that the average symbol energy of the discrete symbol alphabet is one. Finally, $\mathbf{n}\in\mathbb{C}^{N_{r}\times 1}$ is the noise vector with independent and identically distributed (i.i.d.) elements that are distributed according to $\mathcal{CN}(0,\sigma^{2})$, where $\sigma^{2}$ is the noise variance.
III Problem Formulation
We consider a discrete-input continuous-output memoryless channel (DCMC). For equally probable transmitted symbol vectors, the MI is found as [15, 16]
$\text{MI}=\log_{2}(N_{vec})-\frac{1}{N_{vec}}\sum\nolimits_{i=1}^{N_{vec}}\mathbb{E}_{\mathbf{n}}\Bigl{\{}\log_{2}\sum\nolimits_{j=1}^{N_{vec}}e^{\kappa_{i,j}}\Bigr{\}}$ (4)
where $\kappa_{i,j}=(-||\mathbf{H}\mathbf{P}(\mathbf{x}_{i}-\mathbf{x}_{j})+\mathbf{n}||^{2}+||\mathbf{n}||^{2})/\sigma^{2}$. As the direct optimization of the MI in (4) involves discrete variables, which is intractable, we instead consider the optimization of a closed related metric, known as the CR, which is given by
$R_{0}=-\log_{2}\Bigl{[}\frac{1}{N_{vec}^{2}}\sum\nolimits_{i,j=1}^{N_{vec}}e^{-\frac{F_{i,j}}{4\sigma^{2}}}\Bigr{]}$ (5)
where $F_{i,j}=||\mathbf{H}\mathbf{P}(\mathbf{x}_{i}-\mathbf{x}_{j})||^{2}=||\mathbf{H}\mathbf{P}\APLup\mathbf{x}_{i,j}||^{2}$. Since the CR decreases with the argument of the logarithm function in (5), maximizing $R_{0}$ is equivalent to the following optimization problem [16]
$\underset{\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi}}{\operatorname{minimize}}$ $f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})=\sum\nolimits_{i,j=1}^{N_{vec}}e^{-\frac{F_{i,j}}{4\sigma^{2}}}$ (6a)
$\operatorname{subject~{}to}\text{ }$ $\bigl{|}\phi_{n}^{l}\bigr{|}=1,\;\forall l,n$ (6b)
$\bigl{|}\psi_{e}^{k}\bigr{|}=1,\;\forall k,e$ (6c)
$\operatorname{Tr}(\mathbf{P}\mathbf{P}^{H})=N_{s},$ (6d)
where $\boldsymbol{\phi}=[(\boldsymbol{\phi}^{1})^{T},(\boldsymbol{\phi}^{2})^{T},\ldots,(\boldsymbol{\phi}^{L})^{T}]^{T}\in\mathbb{C}^{(NL)\times 1}$, $\boldsymbol{\boldsymbol{\psi}}=[(\boldsymbol{\psi}^{1})^{T},(\boldsymbol{\psi}^{2})^{T},\ldots,(\boldsymbol{\psi}^{K})^{T}]^{T}\in\mathbb{C}^{(MK)\times 1}$, and the equalities in (6b) and (6c) are treated element-wise.
IV Proposed Optimization Method
Input: $\mathbf{P}_{0}$, $\boldsymbol{\phi}_{0}$, $\boldsymbol{\psi}_{0}$, $\nu_{0}>0$, $\mu_{0}^{1:L}>0$, $\tau_{0}^{1:K}>0$, $n\leftarrow 0$, $10<\rho<$ 1, $\delta>0$.
1
2repeat
3
4     repeat /* line search for ${\mathbf{P}}$/
5
6          $\mathbf{P}_{n+1}=\mathcal{P}_{P}(\mathbf{P}_{n}-\nu_{n}\nabla_{\mathbf{P}}f(\mathbf{P}_{n},\boldsymbol{\phi}_{n},\boldsymbol{\psi}_{n}))$
7
8          if $f(\mathbf{P}_{n+1},\boldsymbol{\phi}_{n},\boldsymbol{\psi}_{n})>f(\mathbf{P}_{n},\boldsymbol{\phi}_{n},\boldsymbol{\psi}_{n})-\delta\bigl{\|}\mathbf{P}_{n+1}-\mathbf{P}_{n}\bigr{\|}^{2}$  then
9               $\nu_{n}\leftarrow\rho\nu_{n}$
10                end if
11
12                until  $f(\mathbf{P}_{n+1},\boldsymbol{\phi}_{n},\boldsymbol{\psi}_{n})\leq f(\mathbf{P}_{n},\boldsymbol{\phi}_{n},\boldsymbol{\psi}_{n})-\delta\bigl{\|}\mathbf{P}_{n+1}-\mathbf{P}_{n}\bigr{\|}^{2}$
13               for $l=1$ $\boldsymbol{\mathrm{to}}$ $L$  do
14
15                    repeat /* line search for $\boldsymbol{\phi}^{l}$/
16
17                         $\boldsymbol{\phi}_{n+1}^{l}=\mathcal{P}_{\phi}(\boldsymbol{\phi}_{n}^{l}-\mu_{n}^{l}\nabla_{\boldsymbol{\phi}^{l}}f(\mathbf{P}_{n+1},\bar{\boldsymbol{\phi}}_{n}^{l},\boldsymbol{\psi}_{n}))$
18
19                         if $f(\mathbf{P}_{n+1},\bar{\boldsymbol{\phi}}_{n+1}^{l},\boldsymbol{\psi}_{n})>f(\mathbf{P}_{n+1},\bar{\boldsymbol{\phi}}_{n}^{l},\boldsymbol{\psi}_{n})-\delta\bigl{\|}\boldsymbol{\phi}_{n+1}^{l}-\boldsymbol{\phi}_{n}^{l}\bigr{\|}^{2}$  then
20                              $\mu_{n}^{l}\leftarrow\rho\mu_{n}^{l}$
21                               end if
22
23                               until $f(\mathbf{P}_{n+1},\bar{\boldsymbol{\phi}}_{n+1}^{l},\boldsymbol{\psi}_{n})\leq f(\mathbf{P}_{n+1},\bar{\boldsymbol{\phi}}_{n}^{l},\boldsymbol{\psi}_{n})-\delta\bigl{\|}\boldsymbol{\phi}_{n+1}^{l}-\boldsymbol{\phi}_{n}^{l}\bigr{\|}^{2}$
24                               end for
25
26                              for $k=1$ $\boldsymbol{\mathrm{to}}$ $K$  do
27
28                                   repeat /* line search for $\boldsymbol{\psi}^{k}$/
29
30                                        $\boldsymbol{\psi}_{n+1}^{k}=\mathcal{P}_{\boldsymbol{\psi}}(\boldsymbol{\psi}_{n}^{k}-\tau_{n}^{k}\nabla_{\boldsymbol{\psi}^{k}}(\mathbf{P}_{n+1},\boldsymbol{\phi}_{n+1},\bar{\boldsymbol{\psi}}_{n}^{k})$
31
32                                        if $f(\mathbf{P}_{n+1},\boldsymbol{\phi}_{n+1},\bar{\boldsymbol{\psi}}_{n+1}^{k})>f(\mathbf{P}_{n+1},\boldsymbol{\phi}_{n+1},\bar{\boldsymbol{\psi}}_{n}^{k})-\delta\bigl{\|}\boldsymbol{\psi}_{n+1}^{k}-\boldsymbol{\psi}_{n}^{k}\bigr{\|}^{2}$  then
33                                             $\tau_{n}^{k}\leftarrow\rho\tau_{n}^{k}$
34                                              end if
35
36                                              until $f(\mathbf{P}_{n+1},\boldsymbol{\phi}_{n+1},\bar{\boldsymbol{\psi}}_{n+1}^{k})\leq f(\mathbf{P}_{n+1},\boldsymbol{\phi}_{n+1},\bar{\boldsymbol{\psi}}_{n}^{k})-\delta\bigl{\|}\boldsymbol{\psi}_{n+1}^{k}-\boldsymbol{\psi}_{n}^{k}\bigr{\|}^{2}$
37                                              end for
38
39                                             $n\leftarrow n+1$
40
41                                              until convergence
Algorithm 1 APGM algorithm for solving (6).
We remark that (6) can potentially become a large-scale optimization problem, and thus, first order methods are particularly suitable. Indeed, we adopt the APGM to derive the proposed method for solving (6), which is summarized in 1. Let $(\mathbf{P}_{n},\boldsymbol{\phi}_{n},\boldsymbol{\psi}_{n})$ be the value of $(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ at iteration $n$. Then $\mathbf{P}_{n+1}$ is obtained as
$\mathbf{P}_{n+1}=\mathcal{P}_{P}(\mathbf{P}_{n}-\nu_{n}\nabla_{\mathbf{P}}f(\mathbf{P}_{n},\boldsymbol{\phi}_{n},\boldsymbol{\psi}_{n}))$ (7)
where $\nabla_{\mathbf{P}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ is the gradient of $f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ w.r.t. $\mathbf{P}^{*}$, $\mathcal{P}_{P}(\cdot)$ denotes the projection onto the set defined by (3) and $\nu_{n}$ is the step size, which is found by the line search routine described in steps 1 to 1 of 1.
In optimizing the phase shifts of the SIMs, existing algorithms typically perform a projected gradient step across all phase shifts of all layers simultaneously using the same step size [10, 11]. However, we have observed that such a method results in poor performance (see Fig. 1). To address this, our proposed algorithm optimizes the phase shifts for the transmit and the receive SIMs on a layer-by-layer basis, with each layer being assigned a separate step size. This approach is shown to yield better performance results in our numerical experiments.
More specifically, the phase shifts of the $l$-th layer of the transmit SIM at iteration $n+1$ is determined by
$\boldsymbol{\phi}_{n+1}^{l}=\mathcal{P}_{\boldsymbol{\phi}^{l}}(\boldsymbol{\phi}_{n}^{l}-\mu_{n}^{l}\nabla_{\boldsymbol{\phi}^{l}}f(\mathbf{P}_{n+1},\bar{\boldsymbol{\phi}}_{n}^{l},\boldsymbol{\psi}_{n}))$ (8)
for $l=1,2,\ldots,L$, and where we denote $\bar{\boldsymbol{\phi}}_{n}^{l}=[(\boldsymbol{\phi}_{n+1}^{1})^{T},\ldots,(\boldsymbol{\phi}_{n+1}^{l-1})^{T},(\boldsymbol{\phi}_{n}^{l})^{T},\ldots,(\boldsymbol{\phi}_{n}^{L})^{T}]^{T}$ and $\bar{\boldsymbol{\phi}}_{n+1}^{l}=[(\boldsymbol{\phi}_{n+1}^{1})^{T},\ldots,(\boldsymbol{\phi}_{n+1}^{l})^{T},(\boldsymbol{\phi}_{n}^{l+1})^{T},\ldots,(\boldsymbol{\phi}_{n}^{L})^{T}]^{T}$. In the above, $\nabla_{\boldsymbol{\phi}^{l}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ is the gradient of $f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ w.r.t. $\boldsymbol{\phi}^{l*}$, $\mathcal{P}_{\phi^{l}}(\cdot)$ denotes the projection onto (6b), and $\mu_{n}^{l}$ is the step size for the $l$-th layer. The line search procedure for finding $\mu_{n}^{l}$ is outlined in steps 1 to 1. Similarly, $\boldsymbol{\psi}_{n+1}^{k}$ is found as
$\boldsymbol{\psi}_{n+1}^{k}$ $=\mathcal{P}_{\boldsymbol{\psi}^{k}}(\boldsymbol{\psi}_{n}^{k}-\tau_{n}^{k}\nabla_{\boldsymbol{\psi}^{k}}f(\mathbf{P}_{n+1},\boldsymbol{\phi}_{n+1},\bar{\boldsymbol{\psi}}_{n}^{k})),$ (9)
where $\nabla_{\boldsymbol{\psi}^{k}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ is the gradient of $f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ w.r.t. $\boldsymbol{\psi}^{k*},$ and $\tau_{n}^{k}$ is the step size, which is determined in steps 1 to 1. The involved gradients are provided in 1.
Theorem 1.
The gradients of $f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ with respect to $\mathbf{P}^{*}$, $\boldsymbol{\phi}^{l*}$ and $\boldsymbol{\psi}^{k*}$ are given by
$\nabla_{\mathbf{P}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ $=-\frac{1}{4\sigma^{2}}\mathbf{H}^{H}\mathbf{H}\mathbf{P}\sum\nolimits_{i,j=1}^{N_{vec}}e^{-\frac{F_{i,j}}{4\sigma^{2}}}\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}$ (10)
$\nabla_{\boldsymbol{\phi}^{l}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ $=-\frac{1}{4\sigma^{2}}\sum\nolimits_{i,j=1}^{N_{vec}}e^{-\frac{F_{i,j}}{4\sigma^{2}}}\operatorname{vec}_{d}(\mathbf{L}_{l})$ (11)
$\nabla_{\boldsymbol{\psi}^{k}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ $=-\frac{1}{4\sigma^{2}}\sum\nolimits_{i,j=1}^{N_{vec}}e^{-\frac{F_{i,j}}{4\sigma^{2}}}\operatorname{vec}_{d}(\mathbf{K}_{k})$ (12)
where
$\mathbf{L}_{l}$ $=\boldsymbol{\Theta}^{l+1:L}\mathbf{G}^{H}\mathbf{Z}^{H}\mathbf{H}\mathbf{P}\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}\mathbf{P}^{H}\boldsymbol{\Theta}^{1:l-1}(\mathbf{W}^{l})^{H}$ (13a)
$\mathbf{K}_{k}$ $=(\mathbf{U}^{k})^{H}\boldsymbol{\Upsilon}^{k-1:1}\mathbf{H}\mathbf{P}\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}\mathbf{P}^{H}\mathbf{B}^{H}\mathbf{G}^{H}\boldsymbol{\Upsilon}^{K:k+1}$ (13b)
where $\boldsymbol{\Theta}^{m:n}=(\mathbf{W}^{m})^{H}(\boldsymbol{\mathbf{\Phi}}^{m})^{H}\cdots(\mathbf{W}^{n})^{H}(\boldsymbol{\mathbf{\Phi}}^{n})^{H}$ and $\boldsymbol{\Upsilon}^{m:n}=(\mathbf{\Psi}^{m})^{H}(\mathbf{U}^{m})^{H}\cdots(\mathbf{\Psi}^{m})^{H}(\mathbf{U}^{n})^{H}$.
Proof:
See the Appendix. ∎
After calculating the gradients, the next step is to perform a projection onto the corresponding feasible sets. It is easy to see from (3) that $\mathcal{P}_{P}(\mathbf{P})$ is given by
$\overline{\mathbf{P}}=\mathbf{P}\sqrt{N_{s}/\operatorname{Tr}(\mathbf{P}\mathbf{P}^{H})}.$ (14)
Since the elements of $\boldsymbol{\phi}^{l}$ are constrained to lie on the unit circle, $\mathcal{P}_{\phi}(\boldsymbol{\phi}^{l})$ is defined by
$\overline{\phi_{n}^{l}}=\begin{cases}\phi_{n}^{l}/\bigl{|}\phi_{n}^{l}\bigr{|},&\phi_{n}^{l}\neq 0\\ e^{j\alpha},\alpha\in[0,2\pi]&\phi_{n}^{l}=0.\end{cases}$ (15)
For $\phi_{n}^{l}=0$, the projection can take any point on the unit circle. Finally, $\mathcal{P}_{\boldsymbol{\psi}}(\boldsymbol{\psi}^{k})$ is obtained similarly as in (15). The convergence proof of the proposed method follows the same arguments used in [19] and is omitted for brevity.
V Computational Complexity
In this section, the computational complexity of 1 is analyzed by counting the required number of complex multiplications. We assume that $N\gg N_{t}$ and $E\gg N_{r}$, which is the usual case for a SIM-based HMIMO communication system. We also assume that all matrices $\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}$ are precomputed.
The complexity of the gradient $\nabla_{\mathbf{P}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ is $\mathcal{O}(N_{vec}^{2}N_{s}^{2})$. In addition, $\mathcal{O}(N_{vec}^{2}N_{r}N_{s})$ multiplications are required to obtain $f(\mathbf{P}_{n+1},\boldsymbol{\phi}_{n},\boldsymbol{\psi}_{n})$. Hence, the complexity of computing the precoding matrix $\mathbf{P}$ is equal to $\mathcal{O}(I_{p}N_{vec}^{2}N_{r}N_{s})$, where $I_{p}$ is the number of line search steps.
The complexity of $\sum_{i,j}e^{-\frac{F_{i,j}}{4\sigma^{2}}}\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}$ in (11) is $\mathcal{O}(N_{vec}^{2}N_{s}^{2})$. Furthermore, $\mathcal{O}(LN^{3})$ multiplications are needed to calculate $\mathbf{L}_{l}$. Hence, the complexity of calculating the gradient $\nabla_{\boldsymbol{\phi}^{l}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ is $\mathcal{O}(N_{vec}^{2}N_{s}^{2}+LN^{3})$. After obtaining $\boldsymbol{\phi}_{n+1}^{l}$, the calculation of $\mathbf{B}$ and $\mathbf{H}$ has the complexity of $\mathcal{O}(\text{$N^{3}$}+ENN_{s})$. The complexity of calculating $f(\mathbf{P}_{n+1},\bar{\boldsymbol{\phi}}_{n+1}^{l},\boldsymbol{\psi}_{n})$ is $\mathcal{O}(N_{vec}^{2}N_{r}N_{s})$. Therefore, the complexity of computing the transmit SIM phase shifts $\{\boldsymbol{\phi}^{l}\}_{l=1}^{L}$ can be approximated as $\mathcal{O}(L[LN^{3}+I_{\boldsymbol{\phi}}(\text{$N^{3}$}+ENN_{s}+N_{vec}^{2}N_{r}N_{s})])$, where $I_{\boldsymbol{\phi}}$ is the number of line search loops. Similarly, the complexity of computing the receive SIM phase shifts $\{\boldsymbol{\psi}^{k}\}_{k=1}^{K}$ is equal to $\mathcal{O}(K[KE^{3}+I_{\boldsymbol{\psi}}(\text{$E^{3}$}+ENN_{s}+N_{vec}^{2}N_{r}N_{s})])$, where $I_{\boldsymbol{\psi}}$ is the number of line search steps.
In summary, the complexity of one iteration of the APGM algorithm is given by
$C_{\text{APGM}}=\mathcal{O}(I_{p}N_{vec}^{2}N_{r}N_{s}+L[LN^{3}+I_{\boldsymbol{\phi}}(\text{$N^{3}$}+ENN_{s}$
$+N_{vec}^{2}N_{r}N_{s})]+K[KE^{3}+I_{\boldsymbol{\psi}}(\text{$E^{3}$}+ENN_{s}+N_{vec}^{2}N_{r}N_{s})]).$
VI Simulation Results
In this section, we evaluate the CR and the MI of 1 in a SIM-based HMIMO setup. The channel matrix between the transmit and the receive SIM is modeled based on the spatially-correlated channel model as $\mathbf{G}=\mathbf{R}_{\text{R}}^{1/2}\bar{\mathbf{G}}\mathbf{R}_{\text{T}}^{1/2}\in\mathbb{C}^{E\times L}$ [13, 14] where $\bar{\mathbf{G}}\in\mathbb{C}^{E\times L}$ is distributed according to $\mathcal{CN}(0,\beta\mathbf{I})$; $\beta$ is the free space path loss between the transmit and the receive SIM modeled as $\beta(d)=\beta(d_{0})+10b\log_{10}(d/d_{0})$, where $\beta(d_{0})=20\log_{10}(4\pi d_{0}/\lambda)$ is the free space path loss at the reference distance $d_{0}$, $b$ is the path loss exponent, and $d$ is the distance between the transmitter and the receiver. Moreover, $\mathbf{R}_{\text{T}}\in\mathbb{C}^{L\times L}$ and $\mathbf{R}_{\text{R}}\in\mathbb{C}^{E\times E}$ are the spatial correlation matrices of the transmit and the receive SIM, respectively, and their elements are given by [13].
In the following simulation setup, the parameters are $\lambda=5\,\text{cm}$ (i.e., $f=6\,\text{GHz}$), $N_{t}=2$, $N_{r}=2$, $N_{s}=2$, $\beta=3.5$, $d_{0}=1\,\text{m}$, $d=300\,\text{m}$, $L=K=4$ and $\sigma^{2}=-110\,\text{dB}$. Both the transmit and the receive antennas are placed in arrays parallel to the $x$-axis, and the midpoints of these arrays have coordinates $(0,0,0)$ and $(0,0,d)$, respectively. Also, the inter-antenna separations of these arrays are $\lambda/2$. The meta-atoms in every SIM layer are uniformly placed in a square formation and the size of each meta-atom is $\frac{\lambda}{2}\times\frac{\lambda}{2}$. Moreover, all SIM layers are parallel to the $xy$-plane and their centers are along the $z$-axis. The separation between the neighboring SIM layers, as well as the separation between the first SIM layer and the adjacent antenna array are $\lambda/2$. In the line search procedures in 1, all step sizes are initially set to 1000, $\delta=10^{-3}$ and $\rho=1/2$. The initial values of the optimization variables $\mathbf{P}$, $\boldsymbol{\phi}$, and $\boldsymbol{\psi}$ are randomly generated. All results are averaged over 30 independent channel realizations.
In Fig. 1, we present the CR and the MI of the proposed APGM algorithm for different sizes of the discrete symbol alphabet and different numbers of meta-atoms in SIM layers. In general, the MI reflects the same trend as observed in the CR, albeit always achieving larger values. This justifies the use of the CR as an alternative metric for the MI optimization. In all cases, the CR and the MI reach 90 % of their convergent values in approximately 20 iterations of the proposed algorithm. Also, we notice that the change of $M$ has a negligible influence on the CR, and that the MI may even decrease slightly if $M$ increases when the transmit and the receive SIMs have 49 meta-atoms per layer. This phenomenon is due to the reduced separation between adjacent constellation points when $M$ is increased without a corresponding increase in the average symbol energy, which consequently increases the bit error probability (BEP) per transmission interval. On the other hand, the CR and the MI demonstrate significant improvements as the number of meta-atoms in the SIM layers increases. This enhancement reaffirms the fact that the beamforming capabilities of SIMs are highly dependent on the number of meta-atoms in SIM layers. To show the efficacy of the proposed method, we also present the MI obtained using the same step size for all layers in the SIM. As can be seen clearly, the corresponding MI is not able to converge even after 200 iterations and is significantly lower than the MI obtained using different step sizes for each SIM layer as done in our method.
Figure 1: CR and MI of the proposed APGM algorithm.
Figure 2: MI of the considered system with (w/-) and without (w/o) signal precoding ($M=4$).
In Fig. 2, we show the MI of the considered system with and without signal precoding. It can be clearly observed that signal precoding can substantially increase the MI for about 47 % and 32 % for 49 and 100 meta-atoms per SIM layer, respectively. On the other side, signal precoding alone, i.e., in the absence of the transmit and the receive SIMs, is only able to provide a near zero MI, which is not shown in Fig. 2. Hence, we can conclude that while digital signal precoding of small scale alone shows limited beamforming gain, but its integration with SIM-based HMIMO systems can generate a significant impact on the achievable MI. We also note that the MI is significantly lower when one of the SIMs employs fixed phase shifts, compared to scenarios where both SIMs are tunable. This reduction in the MI is approximately the same regardless of whether the fixed SIM is at the transmit or the receive end.
VII Conclusion
In this paper, we have demonstrated that the MI in a SIM-based HMIMO system can be efficiently optimized with the channel CR. To maximize the CR, we proposed the APGM which optimizes the CR by adjusting the transmit precoding, and the phase shifts for the transmit and the receive SIMs on a layer-by-layer basis. Simulation results showed that the CR is indeed a reliable metric for optimizing the MI in SIM-based HMIMO systems. Also, integrating even a small scale digital precoder in the considered system could substantially increase the MI performance. For future work, an interesting topic is to study the potential of SIMs in the context of emerging integrated sensing and communication systems. Moreover, as a globally optimal solution has not been reported, it remains to be seen how the performance of the proposed method is compared to the optimal performance.
[Proof of 1]
The gradient of $f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ w.r.t. $\mathbf{P}^{*}$ is given by
$\!\!\!\nabla_{\mathbf{P}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})=-\frac{1}{4\sigma^{2}}\sum\nolimits_{i,j=1}^{N_{s}}\exp\left(-\frac{F_{i,j}}{4\sigma^{2}}\right)\nabla_{\mathbf{P}}F_{i,j}.$ (16)
Differentiating $F_{i,j}$, we obtain
$\text{d}F_{i,j}=\operatorname{Tr}\{(\mathbf{H}^{H}\mathbf{H}\mathbf{P}\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H})^{T}\text{d}\mathbf{P}^{*}+\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}\mathbf{P}^{H}\mathbf{H}^{H}\mathbf{H}\text{d}\mathbf{P}\}$
which becomes clear that
$\nabla_{\mathbf{P}}F_{i,j}=\mathbf{H}^{H}\mathbf{H}\mathbf{P}\APLup\mathbf{x}_{i,j}\APLup\mathbf{x}_{i,j}^{H}.$ (17)
Substituting (17) into (16), we obtain (10).
The gradient of $f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})$ w.r.t. $\boldsymbol{\phi}^{l*}$ is given by
$\!\!\!\!\!\!\nabla_{\boldsymbol{\phi}^{l}}f(\mathbf{P},\boldsymbol{\phi},\boldsymbol{\psi})=-\frac{1}{4\sigma^{2}}\sum\nolimits_{i,j=1}^{N_{s}}\exp\left(-\frac{F_{i,j}}{4\sigma^{2}}\right)\nabla_{\boldsymbol{\phi}^{l}}F_{i,j}.$ (18)
Differentiating again $F_{i,j}$ (but now w.r.t. $\boldsymbol{\phi}^{l}$) yields
$\text{d}F_{i,j}$ $=\operatorname{Tr}\{\mathbf{L}_{l}\text{d}(\boldsymbol{\mathbf{\Phi}}^{l})^{H}+\mathbf{L}_{l}^{H}\text{d}\boldsymbol{\mathbf{\Phi}}^{l}\}=\operatorname{Tr}\{\mathbf{L}_{l}^{T}\text{d}\boldsymbol{\mathbf{\Phi}}^{l*}+\mathbf{L}_{l}^{H}\text{d}\boldsymbol{\mathbf{\Phi}}^{l}\}$
$=\operatorname{vec}^{T}(\mathbf{L}_{l})\operatorname{vec}(\boldsymbol{\mathbf{\Phi}}^{l*})+\operatorname{vec}^{T}(\mathbf{L}_{l}^{*})\operatorname{vec}(\boldsymbol{\mathbf{\Phi}}^{l}).$ (19)
Thus, we can conclude that
$\nabla_{\boldsymbol{\phi}^{l}}F_{i,j}=\operatorname{vec}_{d}(\mathbf{L}_{l}),$ (20)
and substituting this gradient into (18), we obtain (11). Following the same steps, we can also prove (12).